OpenStudy (anonymous):

For f(x)= (x-2)^1/5 + 3, calculate the maximum value that delta can be so that f(x) is within 0.2 unit of 3 when x is within delta units of 2.

OpenStudy (blockcolder):

If \(|f(x)-3|\leq0.2\), then \(|(x-2)^{1/5}+3-3|\leq0.2\). If you simplify what's inside the absolute value bars and use the property of absolute value where \(|b^a|=|b|^a\), then you should see the answer.

OpenStudy (anonymous):

why did you put in absolute values signs? and why is there only one side to the inequality.? we've been practiving using 2 sides then choosing the smaller delta or something like that

OpenStudy (anonymous):

@blockcolder could you do \[2.8 \pm (x-2)^{1/5} + 3 < 3.2\] ?

OpenStudy (anonymous):

oops the first sign is supposed to be a less than

OpenStudy (blockcolder):

Yes. The absolute value notation is simply a shorthand. If you use the fact that \[|x-a|\leq b \Leftrightarrow -b\leq x-a \leq b\] Then you can rewrite my absolute value statement as \[\begin{align} |(x-2)^{1/5}+3-3|\leq0.2 &\Leftrightarrow -0.2 \leq (x-2)^{1/5}+3-3\leq 0.2\\ &\Leftrightarrow 2.8 \leq (x-2)^{1/5}+3\leq 3.2 \end{align}\]

OpenStudy (blockcolder):

I thought this absolute value notation was the more commonly taught form.

OpenStudy (anonymous):

maybe we haven't gotten that far yet or something...then you just solve it algebraically, correct? and subtract the smaller delta from the x?

OpenStudy (anonymous):

@blockcolder

OpenStudy (blockcolder):

Lemme just check. The limit definition goes something like this, right? \[|x-a|\leq\delta \Rightarrow |f(x)-L|\leq\epsilon\]

OpenStudy (anonymous):

i believe so

OpenStudy (blockcolder):

Then my absolute value notation is convenient, because if you simplify my expression, then you have \[|(x-2)^{1/5}|\leq0.2 \Rightarrow |x-2|^{1/5}\leq 0.2 \Rightarrow |x-2|\leq 0.2^5\] So that the biggest delta you can use is 0.2^5.

OpenStudy (anonymous):

my books answer is .00032

OpenStudy (blockcolder):

Yeah, and 0.2^5=0.00032 (check with your calculator)

OpenStudy (anonymous):

oh ok thanks!